# zkSTAR: A Zero Knowledge System for Time Series Attack Detection Enforcing Regulatory Compliance in Critical infrastructure Networks

### v2

@@ 1 Introduction
The convergence of information and operational technologies (IT–OT) through widespread IoT adoption has expanded the attack surface for data-driven cyber intrusions. Such attacks often target industrial control systems (ICSs), degrading asset performance and triggering cascading failures across critical infrastructure networks (CINs). While detection occurs locally at utilities, regulatory bodies such as Information Sharing and Analysis Centers (ISACs) aggregate alarms in real time to enhance situational awareness and reduce false positives [dhs_oig2]. The effectiveness of these regulators, however, hinges on the integrity and reliability of local detection data. Recent real-world incidents [langner2011stuxnet, jeffries2022cyber, salazar2024tale] have demonstrated that adversaries can compromise both IT and OT layers to obscure, delay, or disable detection of underlying anomalies. Such incidents expose a fundamental regulatory blind spot pertaining to verifying the integrity of reported detections without re-auditing raw operational data. Moreover, inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions [dhs_oig2]. To address this gap, we introduce zkSTAR, a framework enabling publicly verifiable detection of data-driven ICS attacks in complete zero knowledge of local ICS state-space dynamics.
zkSTAR addresses a critical regulatory gap by enabling provable temporal and statistical consistency verification of ICS detections. Leveraging knowledge-soundness and provably consistent outcomes, it ensures that compromised utilities cannot falsify or suppress alarms even if their IT–OT stacks are breached. The framework enables verifiable regulatory oversight while preserving utility data confidentiality [dhs_oig2], eliminating the need for costly, privacy-intrusive audits that have long impeded secure information sharing [nolan2015cybersecurity]. Additionally, by adopting an on-demand, lazy proof generation paradigm, zkSTAR ensures a low utility overhead as well.
Conventional ICS anomaly detection relies on state-space models that estimate expected sensor behavior and compute residuals to flag deviations from steady-state conditions [li2020detection]. Statistical hypothesis tests on these residuals distinguish normal from abnormal operation, triggering alarms when observed measurements significantly diverge from model predictions [li2020deep]. With existing approaches, regulators can verify such detection outcomes only through full data audits of ICS operations—requiring utilities to relinquish proprietary IoT datasets and raising severe privacy and data-governance concerns. Moreover, utilities must demonstrate both temporal consistency wherein successive state estimates are derived from prior ones; and statistical consistency in which alarms follow valid hypothesis test procedures. These audits are costly, slow, and incompatible with real-time oversight.
@@ 2 Related Works
Attacks on industrial control systems (ICSs) that target IT layers can often be detected through network traffic monitoring [caselli2016specification, ye2004robustness]. However, data-driven attacks that manipulate sensor or control data can evade such approaches by simultaneously impacting both IT and OT layers [urbina2016limiting, kang2014cyber, drias2015taxonomy], potentially causing cascading failures across critical infrastructure. More recently, there have been several attack vulnerabilities that have been discovered such as like Industroyer and Industroyer2 [salazar2024tale], Triton [jeffries2022cyber] and Stuxnet [langner2011stuxnet]. These vulnerabilities have demonstrated the feasibility of joint manipulation of control systems and monitoring data which have motivated the need for robust verifiable detection mechanisms for ICS.
@@
To address some of these challenges, several model-based mechanisms have been proposed [huang2018online, hoehn2016detection, rahman2017multi]. Among these, Kalman filter based approaches are particularly robust and flexible [li2020deep, li2020detection, li2021degradation, ramanan2021blockchain, li2022online], leveraging residuals from state-space models as statistical evidence for anomaly detection. These residual-based tests differentiate routine faults from malicious attacks [li2021degradation] and can support decentralized, network-wide alarm dissemination [ramanan2021blockchain].
@@
From the compliance perspective, however, it is essential for the regulator to receive timely notifications regarding attack detection at utility stakeholders in addition to verifying the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates [ramanan2021blockchain, dhs_oig2]. Therefore, regulatory compliance goals would ideally enable stakeholders to provide publicly verifiable detection outcomes that are accurate and trustworthy without exposing private operational data. Achieving such verifiable compliance under privacy constraints motivates the use of zero-knowledge proofs [narayan2015verifiable], enabling a prover to convince a verifier of correctness without revealing sensitive information.
@@
Originating from foundational work by Goldwasser et al. [goldwasser2019knowledge] and Blum et al. [blum2019non], zero-knowledge proofs evolved into succinct, non-interactive forms (zk-SNARKs) that support efficient verification [kilian1992note, micali2000computationally, gennaro2013quadratic, parno2016pinocchio, ben2013snarks, costello2015geppetto, ben2014succinct]. Recent applications in machine learning verify training integrity and model ownership [huang2022zkmlaas, wang2022ezdps, nguyen2022preserving, yang2023fedzkp]. While existing zk-SNARK frameworks enable general-purpose verifiable computation, none target the verification of dynamical system integrity or hypothesis-test correctness which are cornerstones of industrial attack detection. Recent works on verifiable differential privacy [narayan2015verifiable] and secure multiparty analytics [froelicher2020drynx] focus primarily on data confidentiality but not on verifiable detection correctness, a gap zkSTAR directly addresses. To alleviate this core critical gap, this paper introduces zkSTAR, a framework that enables provable, privacy-preserving verification of Kalman filter–based detection mechanisms, paving the way for regulatory compliance in CINs.
@@ 3 Nonlinear State Space Modeling
For developing zkSTAR, we leverage the extended Kalman Filter (EKF) state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [coskun2017long]. Therefore, we characterize the EKF-based, utility-level attack detection model as a foundational component of zkSTAR for achieving regulatory compliance under zero knowledge. Within zkSTAR, we develop a state-space formulation capable of identifying attack-induced anomalies through locally executed statistical hypothesis tests.
@@ 3.2 Hypothesis Test for Anomaly Detection
We consider residual covariance $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ coupled with residual measurements $r_{t}$ for detecting covert data-driven attacks [li2020detection]. Without loss of generality we note that it is also possible to apply our techniques for detecting false data injection by monitoring $R_{t}$. We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [li2020detection, hwang2017chi]. This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [hwang2017chi].
@@
A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in [li2021degradation]. We define a measure of probabilistic similarity to help establish the distinctness of two sensor measurement vectors as stated in Definition 1.
@@ 4.1 Zero Knowledge Preliminaries
Definition 2 portrays a hashing function that will be utilized to ensure the integrity of state space estimates. We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [menezes1996handbook].
@@ 4.3 Statistical Consistency
Lemma 1 shows that as long as the sensor signal used to carry out EKF predictions is probabilistically similar to the true value of the sensor signal the zkSTAR framework will continue to deliver statistically consistent alarm detection outcomes. Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might in theory go undetected under the zkSTAR framework when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements. Practically, such attacks end up being inconsequential [li2021degradation] since they are unable to effect any material damage to the ICS assets. Therefore, Lemma 1 serves as the basis for Theorem 4.2 which derives an integral link between key invariance, statistical and temporally consistency.
@@ 5.1 Temporal Decomposition Scheme
The temporal decomposition scheme paves the way for zkSNARKs to help enforce integrity of state space predictions based on a pre-committed EKF model. However, the zkSNARK circuit size grows exponentially with the growth in the detection window size making the proof generation step computationally expensive especially for larger systems. Therefore, in order to balance practicality and computational efficiency, we adopt a kernel based strategy that decomposes various compute tasks pertaining to an EKF model. Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to TC and SC attributes of EKF predictions and detection outcomes capable of operating on batched EKF residual outputs. We note that current zkSNARK frameworks such as ezkl lack support for linear solvers, SVD decompositions, and derivative evaluations of complex sub modules due to their high computational cost in zero knowledge [ezkl2024]. To address these limitations, we employ a compute-offline–validate-online strategy, performing such expensive operations off-circuit while verifying their consistency with the state-space model within the zkSNARK circuit.
@@ 5.5 Assumptions and Limitations
The zkSTAR framework relies on several assumptions pertaining to the properties of the zero-knowledge proving system. First, since zkSTAR uses the Halo2 proof generation paradigm as a consequence of using the ezkl framework [ezkl2024], it implicitly depends on elliptic-curve cryptographic primitives and on the integrity of the Common Reference String (CRS). As a result, we assume that the CRS is generated through a publicly verifiable ceremony per-utility. Second, we assume that the EKF model is well-formed and is strongly sensitive to the Jacobian matrices [huang2008analysis] implying that slight deviations in SVD or Kalman gain computations would lead to inconsistent state and residual estimates violating either TC or SC conditions. Third, we assume that verifier is honest-but-curious and that public values of the proof artifacts do not violate privacy of utility state space which is a common requirement of existing regulators [doe]. Finally, we assume that communication channels are assumed to preserve integrity but not confidentiality which can be alleviated by using transport-level encryption like TLS for privacy during proof transmission. Under these assumptions, zkSTAR achieves end-to-end regulatory verification with cryptographic soundness and provable privacy, while maintaining computational feasibility for real-world industrial deployments.
@@ 6 Experimental Results
For experimental studies, we utilize the HAI [hai] and ORNL-PS [pan2015classification, pan2015developing] datasets. To build an asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section 3. The training methodology for EKF models is briefly discussed in Appendix A.2. We encode the temporal and statistical components of these models in corresponding zk-SNARKs as described in previous sections. As a result, the zkSTAR framework encodes temporal and statistical consistency with respect to detection outcomes.
@@
System Implementation Details: All experiments were carried out on a virtual machine (VM) running Ubuntu 24.04 with 100GB of RAM and 16 vCPUs using Python 3.11 with the detection model inference using PyTorch 2.7.1. We utilize the ezkl library [ezkl2024] for generating all the zkSNARKs associated with the zkSTAR framework. We implement zkSTAR using both native and containerized execution modes. The native mode targets high-performance computing (HPC) clusters for evaluating detection accuracy and computational correctness across varying residual aggregation windows ($D$) and precision scale factors (PSF). The containerized mode assesses real-world performance, including CPU and memory usage during witness generation and TC and SC proof generation and verification. Separate containers emulate the utility (prover) and regulator (verifier), communicating via REST APIs and managed through the zkstarctl command-line interface.
@@ 7 Conclusion
Ethics Consideration
None
LLM Usage Considerations
LLMs were used solely for editorial purposes in the preparation of this manuscript (e.g., grammar refinement, stylistic polishing, and clarity improvements). All outputs were manually reviewed, verified, and edited by the authors to ensure technical accuracy and originality. No new ideas, data, or analyses were introduced through the use of LLMs. The literature review, experimental design, results interpretation, and all conceptual contributions were conducted entirely by the authors.
@@ A.1 EKF Modeling for ICS
Our state space modeling framework is based on the extended Kalman filter (EKF) modeling paradigm that is capable of handling non-linearities inherent in the state space process [coskun2017long]. In order to represent the EKF model, we use a deep neural network driven strategy that consists of one feed-forward and three LSTM sub-modules. Specifically, we consider a sensor-driven non-linear system at time $t$, where $x_{t}\in R^{m}$ represents the latent space embedding, $u_{t}\in R^{m}$ represents the control action and $y_{t}\in R^{d}$ represents noisy sensor measurements from asset sensors.
@@
In Equations (23), (24), g,h{g},{h} are the state transition and observation functions respectively. The process and measurement noises at time $t$ are denoted by $v_{t}\in R^{m}$, and $w_{t}\in R^{m}$ respectively. The process and measurement noises follow multivariate normal distributions with zero mean implying that $v_{t}\sim N(0,Q_{t}),w_{t}\sim N(0,R_{t})$, where $Q_{t},R_{t}$ represent the covariance matrices respectively. Such a type of modeling framework has also been used extensively in prior art [smith2011decoupled, mo2014detecting, teixeira2010cyber]. Figure 9 represents the dynamics of our LSTM based deep neural network that is specifically tailored to establish the extended Kalman filter model for ICS subsystems.
@@ B.2 Proof of Lemma 1
We know from Theorem 4.1 that $\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}\text{ , and }\Phi_{t+1}=1$ is a necessary and sufficient condition for temporal consistency. Therefore, to prove this theorem, it is sufficient to show that for a temporally consistent system $\Phi_{t+1}=1\iff\text{Statistical consistency}$. (Case 1 $\Phi_{t+1}=1\implies\text{Statistical consistency}$): Assume that $\rho^{real}_{t+1}\neq\rho^{rep}_{t+1}$ but $\Phi_{t+1}=1$. Based on Lemma 1, we know that there are two probabilistically dissimilar sensor measurement realizations $y^{real}_{t+1},y^{rep}_{t+1}$ corresponding to $\rho^{real}_{t+1},\rho^{rep}_{t+1}$ respectively. This implies that there exists a system of proof artifacts pertaining to $\rho^{rep}_{t+1}$ that satisfy the following relations.
@@
We consider $\mathcal{E}^{\mathcal{U}}$ producing proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in}$. Under the knowledge-soundness property of Halo2 [halo2_zcash_2023], we can state that

### v3

@@ 1 Introduction
The convergence of information and operational technologies (IT–OT) through widespread IoT adoption has expanded the attack surface for data-driven cyber intrusions. Such attacks often target industrial control systems (ICSs), degrading asset performance and triggering cascading failures across critical infrastructure networks (CINs). While detection occurs locally at utilities, regulatory bodies such as Information Sharing and Analysis Centers (ISACs) aggregate alarms in real time to enhance situational awareness and reduce false positives [7]. The effectiveness of these regulators, however, hinges on the integrity and reliability of local detection data. Recent real-world incidents [21, 18, 37] have demonstrated that adversaries can compromise both IT and OT layers to obscure, delay, or disable detection of underlying anomalies. Such incidents expose a fundamental regulatory blind spot pertaining to verifying the integrity of reported detections without re-auditing raw operational data. Moreover, inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions [7]. To address this gap, we introduce zkSTAR, a framework enabling publicly verifiable detection of data-driven ICS attacks in complete zero knowledge of local ICS state-space dynamics.
zkSTAR addresses a critical regulatory gap by enabling provable temporal and statistical consistency verification of ICS detections. Leveraging knowledge-soundness and provably consistent outcomes, it ensures that compromised utilities cannot falsify or suppress alarms even if their IT–OT stacks are breached. The framework enables verifiable regulatory oversight while preserving utility data confidentiality [7], eliminating the need for costly, privacy-intrusive audits that have long impeded secure information sharing [31]. Additionally, by adopting an on-demand, lazy proof generation paradigm, zkSTAR ensures a low utility overhead as well.
Conventional ICS anomaly detection relies on state-space models that estimate expected sensor behavior and compute residuals to flag deviations from steady-state conditions [23]. Statistical hypothesis tests on these residuals distinguish normal from abnormal operation, triggering alarms when observed measurements significantly diverge from model predictions [25]. With existing approaches, regulators can verify such detection outcomes only through full data audits of ICS operations—requiring utilities to relinquish proprietary IoT datasets and raising severe privacy and data-governance concerns. Moreover, utilities must demonstrate both temporal consistency wherein successive state estimates are derived from prior ones; and statistical consistency in which alarms follow valid hypothesis test procedures. These audits are costly, slow, and incompatible with real-time oversight.
@@ 2 Related Works
Attacks on industrial control systems (ICSs) that target IT layers can often be detected through network traffic monitoring [4, 45]. However, data-driven attacks that manipulate sensor or control data can evade such approaches by simultaneously impacting both IT and OT layers [42, 19, 8], potentially causing cascading failures across critical infrastructure. More recently, there have been several attack vulnerabilities that have been discovered such as like Industroyer and Industroyer2 [37], Triton [18] and Stuxnet [21]. These vulnerabilities have demonstrated the feasibility of joint manipulation of control systems and monitoring data which have motivated the need for robust verifiable detection mechanisms for ICS.
@@
To address some of these challenges, several model-based mechanisms have been proposed [16, 13, 35]. Among these, Kalman filter based approaches are particularly robust and flexible [25, 23, 24, 36, 22], leveraging residuals from state-space models as statistical evidence for anomaly detection. These residual-based tests differentiate routine faults from malicious attacks [24] and can support decentralized, network-wide alarm dissemination [36].
@@
From the compliance perspective, however, it is essential for the regulator to receive timely notifications regarding attack detection at utility stakeholders in addition to verifying the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates [36, 7]. Therefore, regulatory compliance goals would ideally enable stakeholders to provide publicly verifiable detection outcomes that are accurate and trustworthy without exposing private operational data. Achieving such verifiable compliance under privacy constraints motivates the use of zero-knowledge proofs [29], enabling a prover to convince a verifier of correctness without revealing sensitive information.
@@
Originating from foundational work by Goldwasser et al. [12] and Blum et al. [3], zero-knowledge proofs evolved into succinct, non-interactive forms (zk-SNARKs) that support efficient verification [20, 27, 11, 34, 1, 6, 2]. Recent applications in machine learning verify training integrity and model ownership [14, 43, 30, 44]. While existing zk-SNARK frameworks enable general-purpose verifiable computation, none target the verification of dynamical system integrity or hypothesis-test correctness which are cornerstones of industrial attack detection. Recent works on verifiable differential privacy [29] and secure multiparty analytics [10] focus primarily on data confidentiality but not on verifiable detection correctness, a gap zkSTAR directly addresses. To alleviate this core critical gap, this paper introduces zkSTAR, a framework that enables provable, privacy-preserving verification of Kalman filter–based detection mechanisms, paving the way for regulatory compliance in CINs.
@@ 3 Nonlinear State Space Modeling
For developing zkSTAR, we leverage the extended Kalman Filter (EKF) state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [5]. Therefore, we characterize the EKF-based, utility-level attack detection model as a foundational component of zkSTAR for achieving regulatory compliance under zero knowledge. Within zkSTAR, we develop a state-space formulation capable of identifying attack-induced anomalies through locally executed statistical hypothesis tests.
@@ 3.2 Hypothesis Test for Anomaly Detection
We consider residual covariance $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ coupled with residual measurements $r_{t}$ for detecting covert data-driven attacks [23]. Without loss of generality we note that it is also possible to apply our techniques for detecting false data injection by monitoring $R_{t}$. We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [23, 17]. This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [17].
@@
A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in [24]. We define a measure of probabilistic similarity to help establish the distinctness of two sensor measurement vectors as stated in Definition 1.
@@ 4.1 Zero Knowledge Preliminaries
Definition 2 portrays a hashing function that will be utilized to ensure the integrity of state space estimates. We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [26].
@@ 4.3 Statistical Consistency
Lemma 1 shows that as long as the sensor signal used to carry out EKF predictions is probabilistically similar to the true value of the sensor signal the zkSTAR framework will continue to deliver statistically consistent alarm detection outcomes. Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might in theory go undetected under the zkSTAR framework when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements. Practically, such attacks end up being inconsequential [24] since they are unable to effect any material damage to the ICS assets. Therefore, Lemma 1 serves as the basis for Theorem 4.2 which derives an integral link between key invariance, statistical and temporally consistency.
@@ 5.1 Temporal Decomposition Scheme
The temporal decomposition scheme paves the way for zkSNARKs to help enforce integrity of state space predictions based on a pre-committed EKF model. However, the zkSNARK circuit size grows exponentially with the growth in the detection window size making the proof generation step computationally expensive especially for larger systems. Therefore, in order to balance practicality and computational efficiency, we adopt a kernel based strategy that decomposes various compute tasks pertaining to an EKF model. Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to TC and SC attributes of EKF predictions and detection outcomes capable of operating on batched EKF residual outputs. We note that current zkSNARK frameworks such as ezkl lack support for linear solvers, SVD decompositions, and derivative evaluations of complex sub modules due to their high computational cost in zero knowledge [40]. To address these limitations, we employ a compute-offline–validate-online strategy, performing such expensive operations off-circuit while verifying their consistency with the state-space model within the zkSNARK circuit.
@@ 5.5 Assumptions and Limitations
The zkSTAR framework relies on several assumptions pertaining to the properties of the zero-knowledge proving system. First, since zkSTAR uses the Halo2 proof generation paradigm as a consequence of using the ezkl framework [40], it implicitly depends on elliptic-curve cryptographic primitives and on the integrity of the Common Reference String (CRS). As a result, we assume that the CRS is generated through a publicly verifiable ceremony per-utility. Second, we assume that the EKF model is well-formed and is strongly sensitive to the Jacobian matrices [15] implying that slight deviations in SVD or Kalman gain computations would lead to inconsistent state and residual estimates violating either TC or SC conditions. Third, we assume that verifier is honest-but-curious and that public values of the proof artifacts do not violate privacy of utility state space which is a common requirement of existing regulators [9]. Finally, we assume that communication channels are assumed to preserve integrity but not confidentiality which can be alleviated by using transport-level encryption like TLS for privacy during proof transmission. Under these assumptions, zkSTAR achieves end-to-end regulatory verification with cryptographic soundness and provable privacy, while maintaining computational feasibility for real-world industrial deployments.
@@ 6 Experimental Results
For experimental studies, we utilize the HAI [38] and ORNL-PS [32, 33] datasets. To build an asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section 3. The training methodology for EKF models is briefly discussed in Appendix A.2. We encode the temporal and statistical components of these models in corresponding zk-SNARKs as described in previous sections. As a result, the zkSTAR framework encodes temporal and statistical consistency with respect to detection outcomes.
@@
System Implementation Details: All experiments were carried out on a virtual machine (VM) running Ubuntu 24.04 with 100GB of RAM and 16 vCPUs using Python 3.11 with the detection model inference using PyTorch 2.7.1. We utilize the ezkl library [40] for generating all the zkSNARKs associated with the zkSTAR framework. We implement zkSTAR using both native and containerized execution modes. The native mode targets high-performance computing (HPC) clusters for evaluating detection accuracy and computational correctness across varying residual aggregation windows ($D$) and precision scale factors (PSF). The containerized mode assesses real-world performance, including CPU and memory usage during witness generation and TC and SC proof generation and verification. Separate containers emulate the utility (prover) and regulator (verifier), communicating via REST APIs and managed through the zkstarctl command-line interface.
@@ A.1 EKF Modeling for ICS
Our state space modeling framework is based on the extended Kalman filter (EKF) modeling paradigm that is capable of handling non-linearities inherent in the state space process [5]. In order to represent the EKF model, we use a deep neural network driven strategy that consists of one feed-forward and three LSTM sub-modules. Specifically, we consider a sensor-driven non-linear system at time $t$, where $x_{t}\in R^{m}$ represents the latent space embedding, $u_{t}\in R^{m}$ represents the control action and $y_{t}\in R^{d}$ represents noisy sensor measurements from asset sensors.
@@
In Equations (23), (24), g,h{g},{h} are the state transition and observation functions respectively. The process and measurement noises at time $t$ are denoted by $v_{t}\in R^{m}$, and $w_{t}\in R^{m}$ respectively. The process and measurement noises follow multivariate normal distributions with zero mean implying that $v_{t}\sim N(0,Q_{t}),w_{t}\sim N(0,R_{t})$, where $Q_{t},R_{t}$ represent the covariance matrices respectively. Such a type of modeling framework has also been used extensively in prior art [39, 28, 41]. Figure 9 represents the dynamics of our LSTM based deep neural network that is specifically tailored to establish the extended Kalman filter model for ICS subsystems.
@@ B.2 Proof of Lemma 1
We know from Theorem 4.1 that $\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}\text{ , and }\Phi_{t+1}=1$ is a necessary and sufficient condition for temporal consistency. Therefore, to prove this theorem, it is sufficient to show that for a temporally consistent system $\Phi_{t+1}=1\iff\text{Statistical consistency}$.
(Case 1 $\Phi_{t+1}=1\implies\text{Statistical consistency}$): Assume that $\rho^{real}_{t+1}\neq\rho^{rep}_{t+1}$ but $\Phi_{t+1}=1$. Based on Lemma 1, we know that there are two probabilistically dissimilar sensor measurement realizations $y^{real}_{t+1},y^{rep}_{t+1}$ corresponding to $\rho^{real}_{t+1},\rho^{rep}_{t+1}$ respectively. This implies that there exists a system of proof artifacts pertaining to $\rho^{rep}_{t+1}$ that satisfy the following relations.
@@
For detection suppression, we assume that the compromised utility possesses an extractor $\mathcal{E}^{\mathcal{U}}$ that produces a valid proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in},\Delta_{t}^{out})\neq\Omega_{t}$, with a non-negligible probability. We consider $\mathcal{E}^{\mathcal{U}}$ producing proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in},\Delta_{t}^{out})$. Under the knowledge-soundness property of Halo2 [46], we can say